Joint ABT/BI Technical Meeting

865/2-C19 (CERN)



Show room on map
    • 10:30 10:45
      Follow-up of open actions 15m
      • TPSC4.2 spare: BI system installation moved to Q3/2020.
      • Morad to contact Cedric Baud concerning follow up of BBSH22778.
      • Plan repair of TPS15 BI system removed from PS during LS2 (Ray).
      • TPS15 installation in PS to be completed (camera and cover modification) (Ray).
    • 10:45 10:50
      Outlook for ABT needs from BI in next 6-12 months 5m
    • 10:50 11:00
      Review of ABT/BI sharepoint list 10m
    • 11:00 11:10
      AOB and round table 10m